# The impact of the eccentricity on the collapse of an ellipsoid into a black hole

### v1

@@ III The black hole formation
Then we find the moment when $R=Z$. The spheroid occupies the smallest possible volume at this moment and therefore has the highest density inside a sphere. We have all the reasons to consider the value of $R(t)$ at this time as a better estimation of $R_{fin}$, which we call ’$R_{fin}$ with flyby’.
If we consider a prolate spheroid, we should find ’$Z_{fin}$ with flyby’ instead, of course. To do so, we act exactly as in the previous case, but we reverse the $R$ component of the velocity $v_{R}\rightarrow-v_{R}$ instead of $v_{Z}$ and substitute to (15) the equations for $\alpha$ and $\beta$ corresponding to the prolate case (see (8, 9)). Then we find the moment after $t_{col}$ when $R=Z$ and name the value of $Z(t)$ at this moment ’$Z_{fin}$ with flyby’.
@@ IV Discussion
Let us apply the obtained results. Hereafter, we assume the standard $\Lambda$CDM cosmology with the present-day radiation fraction $\Omega_{m,0}=5.0\times 10^{-5}$ and $H_{0}=71$ (km/s) Mpc-1 [6]. Then the present-day critical density is $\rho_{c,0}\approx 0.97\cdot 10^{-29}$ g/cm3.

### v2

@@ III The black hole formation
Then we find the moment when $R=Z$. The spheroid occupies the smallest possible volume at this moment and therefore has the highest density inside a sphere. We have all the reasons to consider the value of $R(t)$ at this time as a better estimation of $R_{fin}$, which we call $R_{fin}$ with flyby’.
If we consider a prolate spheroid, we should find $Z_{fin}$ with flyby’ instead, of course. To do so, we act exactly as in the previous case, but we reverse the $R$ component of the velocity $v_{R}\rightarrow-v_{R}$ instead of $v_{Z}$ and substitute to (15) the equations for $\alpha$ and $\beta$ corresponding to the prolate case (see (8, 9)). Then we find the moment after $t_{col}$ when $R=Z$ and name the value of $Z(t)$ at this moment $Z_{fin}$ with flyby’.
@@ IV Discussion
[…]